OpenStudy (anonymous):

Find the limit as arctan(2n/ (2n+1)) when n approaches infinity.

OpenStudy (irishboy123):

?

OpenStudy (anonymous):

\[\lim_{2n\to \infty}\frac{2n}{2n+1}=1\]

OpenStudy (anonymous):

But it is arctan?

OpenStudy (anonymous):

yes and artan is a continuous function

OpenStudy (anonymous):

I'm sorry, but how does the relate? :)

OpenStudy (anonymous):

that*

OpenStudy (anonymous):

making your answer \[\arctan(1)\]

OpenStudy (anonymous):

So we are able to find the limit of the inner function then apply the outer function to the result of the inner limit?

OpenStudy (anonymous):

yes, if it is continuous